A corporate bond specifies that interest of $400 is to be paid every six months for 20 years, and then the $1,000 principal or face value of the bond also will be paid. The market interest rate for bonds of this quality is 10 percent per year. What is the market value of this bond?
1
Expert's answer
2015-03-03T09:59:33-0500
If a corporate bond specifies that interest of $400 is to be paid every six months for 20 years, and then the $1,000 principal or face value of
the bond also will be paid, and if the market interest rate for bonds of
this quality is 10 percent per year. The market value of this bond will
be:
P = C*((1 - (1 + i)^-n)/i) + M*(1 + i)^-n,
where:
F = face values
C = coupon payment
N = number of payments
i = market interest rate
M = face value
P = market price of bond.
P = 400*((1 - (1 + 0.05)^(-20))/0.05) + 1000*1.05^(-20) = $5361.77